Time of last reading:
March 20, 2019, 3 p.m.
Jan. 23, 2018, 10:15 a.m.
Issue Feb. 12, 2018 - Feb. 22, 2018
It has been brought to our attention by Environmental Instruments Ltd (manufacturers of AQMesh) that due to a bug in a server systems update that NO2 and O3 data from all AQMesh pods were effected from 7am on 12th February until 9am on 22nd February 2018. No other channels are affected. We caution you against using NO2 and O3 data for this period until the corrections have been applied. We urge users to re-call this data from Tuesday 27th when all readings will have been re-processed. The Urban Observatory apologises for the inconvenience.Issue April 6, 2018 - April 6, 2018
NO data reading negative. Sensor needs replacing. NO2 data reading too high. Sensor needs replacing.Issue April 9, 2018 - April 9, 2018
Replaced faulty N02 sensor at Corner House AQ MeshIssue May 1, 2018 - May 1, 2018
Aq_meshes 1904 (Chillingham Rd Primary School) , 1905 (Corner House, Heaton) removed for slope/offset adjustment at Precision SiteIssue May 22, 2018 - May 22, 2018
Deployed Coast Road Corner House aq_mesh1971150 as replacement for aq_mesh1905150 Lamp post 8S235 Post Code NE6 5QD GPS 54.99113938 -1.583788097Issue May 22, 2018 - May 22, 2018
Chillingham Road Primary (1904150) and Coast Road Corner House (1095150) were not 'refurbished' back in Jan 18 with the rest of that batch. These have been removed and substituted for new AQ Mesh pods ie Chillingham Road Primary is now aq_mesh1970150 Coast Road Corner House is now aq_mesh1971150 . Please refer to those pods from this date till further notice. ThanksIssue Aug. 2, 2018 - Aug. 2, 2018
Redeployed ( formerly SenseMyStreets Coast Rd/Corner House) Now at St Teresa's Primary School Heaton Rd Lamppost 27H116 NE6 5HN 54.984003 -1.58199Issue Sept. 7, 2018 - Sept. 7, 2018
Relocated to Gold Node aq_mesh1968150 @ Tyne Bridge Precision Station approx 13:30Issue Oct. 3, 2018 - Oct. 3, 2018
aq_mesh1905150 ST TERESA'S PRIMARY SCHOOL HEATON Started 07/09/2018 at TYNE_BRIDGE 02/10/18 Linear Regression applied from 07/09/18 03/10/18 Re-instated at original site Precision Station TYNE_BRIDGE 1056100 (from 08/09/18 to 01/10/2018) 1905150 CO (co-location data from 08/09/2018 to 15/09/2018) Original slope:0.734700024 Original Offset:180.7640076 New Slope:0.00096202 New Offset:1.021187404 R square:0.600313234 NO Original slope:1 Original Offset:-1.870200038 New Slope:0.895772732 New Offset:1.513792243 R square:0.949827668 NO2 Original slope:1 Original Offset:7.512499809 New Slope:0.992272999 New Offset:4.256128404 R square:0.8676163 O3 Original slope:0.598399997 Original Offset:9.694499969 New Slope:0.348258852 New Offset:13.74364994 R square:0.881707131 PM10 Original slope:1 Original Offset:0 New Slope:2.293259122 New Offset:2.401365391 R square:0.544444393 PM2.5 Original slope:1 Original Offset:0 New Slope:4.048410298 New Offset:0.456312641 R square:0.793511745 PM1 Original slope:1 Original Offset:0 New Slope:7.356609394 New Offset:0.202757435 R square:0.823239657Issue March 4, 2018 - April 10, 2018
NO2 Sensor failed 04/03/18. Replaced 10/04/18Issue Oct. 24, 2018 - Oct. 25, 2018
All AQ Meshes did not report any data between Weds 24th Oct 06:00 – Thurs 25th Oct 11:15, reason not yet knownIssue Nov. 16, 2018 - Nov. 16, 2018
Approx 15:00 - 16:00 : Changed the sampling frequency for gases for all AQ Meshes to be standard across all pods : P1 (Gas Reading Frequency) - Ten secs (unchanged) P2 (Gas Averaging Frequency) - One Minute (was mostly 15 mins) P3 (Gas Transmission Frequency) - 30 Mins (was One Hour)Issue Nov. 22, 2018 - Nov. 23, 2018
22/11/18 17:00 AirMonitors carrying out server maintenance at approx 17:00 (UTC) on Thursday November 22nd 23/11/10:00 server back. Possible data loss during this periodIssue Dec. 6, 2018 - Dec. 6, 2018
06/12/2018 Major disruption to the 02 Cellphone/Data network today. This will affect many sensors that rely on that network (eg Emotes)Issue Jan. 20, 2019 - Jan. 21, 2019
20/01/19 17:00 approx Sensors data not being displayed on the portal. This was due to failed video data stuck in the queue and blocking all other data.Issue Jan. 15, 2019 - Jan. 16, 2019
15/01/19 02:15 - 16/01/19 Data lost from all sensors due to quue failure. This data may eventually get back-filled.
Sensor Maintenance Records: